OpenStudy (anonymous):

What is the directrix of a parabola whose equation is (y + 3)2 = 8(x − 3)?

OpenStudy (jdoe0001):

\(\large { \begin{array}{llll} (y-{\color{blue}{ k}})^2=4{\color{purple}{ p}}(x-{\color{brown}{ h}}) \\ \end{array} \qquad \begin{array}{llll} vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\\ {\color{purple}{ p}}=\textit{distance from vertex to }\\ \qquad \textit{ focus or directrix} \end{array} \\ \quad \\ (y + 3)^2 = 8(x - 3) }\) any ideas on the vertex coordinates? what about theh "p" distance?

OpenStudy (mathmate):

The directrix of a parabola y^2=4px is x=-p Translating by the vertex by (3,-3), the parabola becomes 8(x-3)=(y+3)^2 from which we can equate 4p=8, or p=2 units to the left of the vertex, which means that the directrix is at x=3-2=1.

OpenStudy (mathmate):